# Dynamic Knowledge Integration for Evidence-Driven Counter-Argument Generation with Large Language Models

### v1

@@ Abstract
This paper investigates the role of dynamic external knowledge integration in improving counter-argument generation using Large Language Models (LLMs). While LLMs have shown promise in argumentative tasks, their tendency to generate lengthy, potentially unfactual responses highlights the need for more controlled and evidence-based approaches. We introduce a new manually curated dataset of argument and counter-argument pairs specifically designed to balance argumentative complexity with evaluative feasibility. We also propose a new LLM-as-a-Judge evaluation methodology that shows a stronger correlation with human judgments compared to traditional reference-based metrics. Our experimental results demonstrate that integrating dynamic external knowledge from the web significantly improves the quality of generated counter-arguments, particularly in terms of relatedness, persuasiveness, and factuality. The findings suggest that combining LLMs with real-time external knowledge retrieval offers a promising direction for developing more effective and reliable counter-argumentation systems. Data and code publicly available. [^1] https://anonymous.4open.science/r/counter-argument-generation/
@@ 1 Introduction
It is possible to distinguish two main research lines in Argumentation in NLP: (i) argument mining, which involves analyzing unstructured texts to automatically identify and extract argumentative elements Cabrio and Villata (2018); Stab and Gurevych (2017); Yeginbergen et al. (2024); (ii) argument generation, which focuses on generating argumentative texts using external knowledge sources Hua et al. (2019); Schiller et al. (2021) or summarizing key argumentative points Roush and Balaji (2020); Syed et al. (2021); Chen et al. (2024).
@@
This paper investigates whether dynamic integration of external knowledge helps Large Language Models (LLMs) to improve counter-argument generation. Counter-argument generation seeks to develop an effective framework for presenting alternative perspectives to an argument while ensuring the correctness of the message and incorporating factual evidence Wachsmuth et al. (2017, 2018). LLMs have shown promising potential to deal with debates or any disagreements by solely relying on the parametric knowledge encoded in the model Chen et al. (2024); Alshomary and Wachsmuth (2023). Moreover, most of the LLMs have security safeguarding to avoid harmful interactions by any means Bai et al. (2022); Ouyang et al. (2022); Touvron et al. (2023). While maintaining the harmlessness is recommendable, it is also important to be reasonable, persuasive, and grounded with a factual basis when arguing Khan et al. (2024).
@@
Previous work on counter-argument generation has focused on specific aspects of the argument Schiller et al. (2021); Saha and Srihari (2023), integrating different external news sources Hua et al. (2019), or by attacking different evidences of the initial argument Jo et al. (2020); Chen et al. (2024). However, no previous work has considered integrating dynamic knowledge to inform counter-argument generation. Furthermore, we argue that existing datasets for counter-argument generation consist of too long Hua and Wang (2018); Hua et al. (2019); Chen et al. (2024) or too short Schiller et al. (2021); Lin et al. (2023) counter-arguments which makes it difficult to accurately evaluate their quality. Thus, although modern LLMs tend to generate rather lengthy essay-style responses that may be highly persuasive but lack coherence, logic, and factual evidence, restricting them to short single sentences is insufficient to study the complexity and desired pragmatic characteristics of a good counter-argument. Finally, the manual evaluation of counter-argument quality is a challenging, time-consuming, and highly subjective task that requires knowledge of the subject matter and credible evidence to support or refute it Wachsmuth et al. (2017); Wang et al. (2017); Hua et al. (2019). In this sense, we believe that traditionally used Alshomary et al. (2021); Chen et al. (2024) reference-based automatic metrics such as BLEU, METEOR or BERTScore fail to accurately capture the nuanced relationship between the generated counter-argument and the human judgment.
@@
In order to address these issues, we propose using retrieved external knowledge from dynamic sources without any limits to particular outlets or databases, along with LLMs, to generate efficient, concise evidence-based counter-arguments. Furthermore, we present a new curated dataset of concise and structured human-generated argument and counter-argument pairs in which the length of the counter-arguments is enough to study the main argumentative aspects while facilitating manual and automatic evaluation. While this paper employs both manual and automated evaluation methods, we introduce a novel automated evaluation approach using LLMs-as-a-Judge, designed to optimize correlation with human assessments.
Using the new dataset and dynamically retrieved external evidence, this work aims to answer the following research questions (RQ): RQ1: Does the integration of dynamic external knowledge into LLMs help to generate better counter-argumentation? RQ2: Which automatic evaluation method correlates better with human judgments? RQ3: To what extent do LLMs use retrieved external evidence in producing counter-arguments?
@@
Figure 1 illustrates our proposed framework. First, we automatically generate queries (in the form of critical questions) that challenge the main points of the argument or claim and feed those queries to the web search. Next, related evidence is retrieved, and, lastly, the claim and the retrieved evidence are provided as context to the LLM to generate a counter-argument.
@@
The main contributions of our paper are the following: (i) we publicly share a new dataset of manually curated arguments and counter-arguments; (ii) we introduce a new method to dynamically integrate external knowledge retrieved from the web in LLM-based counter-argument generation; (iii) experimental results show that the generation of counter-arguments with LLMs is improved through the integration of dynamic external knowledge, with factual evidence demonstrating a particularly significant impact on pragmatic aspects including relatedness, persuasiveness, and factuality and (iv), our automatic evaluation based on LLM-as-a-Judge reveals a higher correlation with human judgments compared with reference-based metrics such as BLEU, METEOR, or BERTscore.
@@ 2 Related Work
Prior research in efficient and persuasive argument generation has approached the problem from various perspectives focusing on different aspects of arguments. For instance, Jo et al. (2020) investigated the ability of machine learning approaches to detect attackable sentences in the arguments, and they concluded that automatic approaches are more stable in this task than human annotators depending on the subjectivity, topic, and tone of the argument. Alshomary et al. (2021) focused on generating counter-arguments by pinpointing and challenging weak premises. Schiller et al. (2021) produce arguments by specifying the desired aspect and stance in a sentence-level setting via Conditional Transformer Language Model (CTRL) (Keskar et al., 2019). Similarly, Saha and Srihari (2023) proposed a method to control both the topic and stance of an argument while enriching it with factual evidence using an encoder-decoder language model. Alshomary and Wachsmuth (2023) propose a counter-argument generation based on the stance of the conclusion of the argument. Lin et al. (2023) employed large language models for sentence-level counter-argument generation, implementing a Chain-of-Thought methodology. Finally, Chen et al. (2024) evaluated LLMs in several argument mining and generation tasks showing the potentiality of LLMs for this particular task.
The use of external sources has demonstrated its effectiveness in the generation of alternative perspectives. Wachsmuth et al. (2018) analyzed the question of retrieving the best counter-arguments when no prior knowledge is persistent. Hua and Wang (2018) introduced a framework that incorporated information retrieval, but their approach was limited to using the Wikipedia database as the external source. However, Wikipedia primarily contains static factual information, which may not align with the dynamic nature of arguments. To address this limitation, Stab et al. (2018) expanded the scope by indexing all documents from the Common Crawl database for argument retrieval. Hua et al. (2019) proposed an enhanced framework that leverages databases from news outlets alongside Wikipedia to retrieve evidence and improve the quality of the generated counter-arguments.
@@ 3 Data
In order to perform our experiments and avoid input and output length inequality in the data, we constructed a new corpus for the evaluation of counter-argument generation. Previous work often focused on either (too short) sentence-level (Lin et al., 2023; Schiller et al., 2021) or (too long) paragraph-level (Hua et al., 2019; Hua and Wang, 2018) generation.
Given the capabilities of modern LLMs, it was essential to define clear input and output data to ensure accurate, robust, and fair comparisons during evaluation. Thus, our objective is to create a dataset of argument/counter-argument pairs that would meet specific criteria. We argue that (i) generating single-sentence claim-based counter-arguments is insufficient to accurately assess the quality of counter-arguments produced by LLMs and (ii), too long counter-arguments make it extremely difficult to properly evaluate the argumentative quality of the generated text. Our analysis revealed that Large Language Models (LLMs), when not given explicit length constraints, tend to generate verbose, essay-like responses that frequently deviate from true argumentative form, lack substantiating evidence, and demonstrate poor coherence with the original argument while overemphasizing persuasive elements. To address these limitations, we propose generating counter-arguments with a maximum length of three sentences, focusing on conciseness, factual content, and direct alignment with the input argument.
With this aim in mind, we constructed a new dataset of argument and counter-argument pairs using the CANDELA corpus as a basis Hua et al. (2019). The corpus consists of debates and discussions on various controversial topics from the r/ChangeMyView [^2] https://www.reddit.com/r/changemyview subreddit. The corpus is centered around real-world online debates where users post their opinions and evidence for any controversial topic and expect other users to provide reasoning for an alternative perspective.
@@
Original 16 1996
Intermediate 3 511
Final 3 387
@@
Original 30 4478
Intermediate 5 845
Final 3 460
@@
CANDELA is available in a format split by sentenced, tokenized, and lowercase, making the reconstruction of the corpus necessary. To address this, we employed LLMs to convert the data back into a coherent, human-readable format.
Once the data was fully reconstructed in a human-readable format, we summarized all arguments and counter-arguments. To avoid any bias we choose to use a language model different from those used in our experimentation, namely, Llama-3.1-70B instruct Dubey et al. (2024). All summaries were then manually verified against the original data and re-summarized where needed to ensure the semantic and pragmatic correctness of the content.
While the corpus includes data from real-world interactions and reflects arguments from natural exchanges of information, we found that not all topics were suitable for counter-argument generation with external knowledge. Specifically, topics lacking a scientific or factual foundation, such as philosophical questions or deeply subjective topics, often trigger LLMs to produce generic, safe responses due to their safety guardrails.
@@
Finally, we further manually refined the summaries to follow a structured argumentative format, emphasizing components such as the main claim, supporting evidence, and examples where applicable. This process resulted in a dataset of 150 high-quality 3-sentence paragraph-level argument and counter-argument pairs. The final version of the corpus used in the experiments can be seen in the Appendix 3. The average distribution of sentences and tokens in the data is shown in Table 1.
@@ 4 Experimental setup
External knowledge is obtained through the web search provided by the Cohere API [^3] https://cohere.com/. The process involved automatically generating five queries (averaging 87 words each) designed to challenge the validity and veracity of the original argument’s claims and premises by questioning key points that required factual substantiation. These queries were sequentially submitted to the web search engine, and the retrieved results, averaging 5,496 words in length, were incorporated as contextual information in the final prompt presented to the language model. The model then generated counter-arguments based on both the original argument and the retrieved contextual information.
@@
This experimentation is performed with two instruct LLMs with strong results on text generation tasks in two different configurations: (i) the LLM using only the claim and its parametric knowledge as prompt and (ii) the LLM with dynamic external knowledge. This results in the following four models:
@@
Importantly, all experiments are conducted in a zero-shot inferential setting to assess the real capabilities of LLMs in counter-argument generation. This setup ensures a fair and robust evaluation of their performance in generating meaningful, well-reasoned, and factual counter-arguments.
@@ 4.1 Evaluation
It should be noted that the evaluation of generated counter-arguments is inherently subjective. While human evaluation is the gold standard, it is time-intensive, costly, and prone to individual biases Wachsmuth et al. (2017); Hua et al. (2019); Chen et al. (2024). To address these limitations, along with human evaluation, we will also provide two types of automatic evaluation. First, using reference-based metrics such as BLEU, METEOR or BERTscore previously used in counter-argumentation generation Alshomary et al. (2021); Chen et al. (2024). Second, we propose to use the LLM-as-a-Judge approach. Leveraging the consistency, scalability, and efficiency of the LLMs, this method enables rapid and reproducible scoring across the five dimensions. To the best of our knowledge, we are the first to propose an LLM-based evaluation for counter-argument generation.
@@
Human evaluation. We recruited human evaluators through the crowdsourcing platform Prolific [^4] https://www.prolific.com/ to assess the quality of a sample of 75 generated counter-arguments across the five predefined dimensions. We performed two rounds of evaluations of three participants to ensure that the obtained manual judgments were of high quality. We set test questions in the questionnaire to determine whether the evaluations were performed fairly. Subsequently, we selected the four participants who correctly passed the test. All evaluators were compensated above the minimum rate recommended by the platform. An example of instructions provided for manual evaluation can be found in Appendix C.
@@
Reference-based evaluation. Following the previous evaluations on counter-argument generation Hua et al. (2019); Lin et al. (2023); Chen et al. (2024) we evaluate using reference-based overlap and similarity metrics, such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), METEOR (Banerjee and Lavie, 2005) and BERTScore (Zhang et al., 2019). BLEU and ROUGE compare the overlap between the counter-argument and the claim whereas METEOR considers also synonyms, paraphrases, and word stems. Finally, BERTScore takes into account the semantic context and meaning of the text, going beyond surface-level word matching.
Ranking. Based on the evaluation scores per dimension, we obtain rankings of each the 5 candidate counter-arguments (4 automatically generated and the gold reference) by summing the scores for each counter-argument, as provided by the human and LLM-as-a-Judge evaluators. Formally, let the evaluation process involve $n$ counter-arguments, $m$ dimensions, and $e$ evaluators. Let $Score_{i,j,d}$ represent the score assigned by evaluator $j$ for counter-argument $i$ on dimension $d$.

### v2

@@ Abstract
This paper investigates the role of dynamic external knowledge integration in improving counter-argument generation using Large Language Models (LLMs). While LLMs have shown promise in argumentative tasks, their tendency to generate lengthy, potentially non-factual responses highlights the need for more controlled and evidence-based approaches. We introduce a reconstructed and manually curated dataset of argument and counter-argument pairs specifically designed to balance argumentative complexity with evaluative feasibility. We also propose a new LLM-as-a-Judge evaluation methodology that shows a stronger correlation with human judgments compared to traditional reference-based metrics. Our experimental results demonstrate that integrating dynamic external knowledge from the web significantly improves the quality of generated counter-arguments, particularly in terms of relatedness, persuasiveness, and factuality. The findings suggest that combining LLMs with real-time external knowledge retrieval offers a promising direction for developing more effective and reliable counter-argumentation systems. Data and code are publicly available. [^1] https://github.com/anaryegen/counter-argument-generation
@@ 1 Introduction
It is possible to distinguish two main research lines in Argumentation in NLP: (i) argument mining, which involves analyzing unstructured texts to automatically identify and extract argumentative elements and their relations Cabrio and Villata (2018); Stab and Gurevych (2017); Yeginbergen et al. (2024); (ii) argument generation, which focuses on generating argumentative texts using external knowledge sources Hua et al. (2019); Schiller et al. (2021) or summarizing key argumentative points Roush and Balaji (2020); Syed et al. (2021); Chen et al. (2024).
@@
This paper contributes to the second line of research by investigating whether dynamic integration of external knowledge helps LLMs to improve counter-argument generation. Counter-argument generation seeks to develop an effective framework for presenting alternative perspectives to an argument while ensuring the correctness of the message and incorporating factual evidence Wachsmuth et al. (2017, 2018). LLMs have shown promising potential to deal with debates or any disagreements by solely relying on the parametric knowledge encoded in the model Chen et al. (2024); Alshomary and Wachsmuth (2023). Moreover, most of the LLMs have security safeguards to avoid harmful interactions by any means Bai et al. (2022); Ouyang et al. (2022); Touvron et al. (2023). While maintaining harmlessness is recommendable, it is also important to be reasonable, persuasive, and grounded with a factual basis when arguing Khan et al. (2024).
@@
Previous work on counter-argument generation has focused on specific aspects of the argument Schiller et al. (2021); Saha and Srihari (2023), integrating different external news sources Hua et al. (2019), or by attacking different evidence of the initial argument Jo et al. (2020); Chen et al. (2024). However, no previous work has considered integrating dynamic knowledge to inform counter-argument generation. Furthermore, we argue that existing datasets for counter-argument generation consist of excessively long Hua and Wang (2018); Hua et al. (2019); Chen et al. (2024) or extremely short Schiller et al. (2021); Lin et al. (2023) arguments, which makes it difficult to accurately evaluate their quality. Thus, although modern LLMs tend to generate rather lengthy essay-style responses that may be highly persuasive, they lack coherence, logic, and factual evidence, restricting them to short single sentences is insufficient to study the complexity and desired pragmatic characteristics of a good counter-argument Hinton and Wagemans (2023); Ji et al. (2023); Carrasco-Farre (2024); Verma et al. (2024). Finally, the manual evaluation of counter-argument quality is a challenging, time-consuming, and highly subjective task that requires knowledge of the subject matter and credible evidence to support or refute it Wachsmuth et al. (2017); Wang et al. (2017); Hua et al. (2019). In this sense, we believe that traditionally used Alshomary et al. (2021); Chen et al. (2024) reference-based automatic metrics such as BLEU, METEOR, or BERTScore fail to accurately capture the nuanced relationship between the generated counter-argument and the human judgment.
@@
In order to address these issues, we propose using retrieved external knowledge from dynamic sources without any limits to particular outlets or databases, along with LLMs, to generate efficient, concise, evidence-based counter-arguments. Furthermore, we present a new curated dataset of concise and structured human-generated argument and counter-argument pairs in which the length of the counter-arguments is enough to study the main argumentative aspects while facilitating manual and automatic evaluation. While this paper employs both manual and automated evaluation methods, we introduce a novel automated evaluation approach using LLMs-as-a-Judge, designed to optimize correlation with human assessments.
Using the new dataset and dynamically retrieved external evidence, this work aims to answer the following research questions (RQ):
  • RQ1: Does the integration of dynamic external knowledge into LLMs help to generate better counter-argumentation?
  • RQ2: Which automatic evaluation method correlates better with human judgments?
  • RQ3: To what extent do LLMs use retrieved external evidence in producing counter-arguments?
@@
Figure 1 illustrates our proposed framework. First, we automatically generate queries that challenge the main points of the argument or claim for more fact-based information and feed those queries to the web search. Next, related evidence is retrieved, and, lastly, the claim and the retrieved evidence are provided as context to the LLM to generate a counter-argument. The main contributions of our paper are the following:
  1. 1.
    We publicly share a new dataset of manually curated arguments and counter-arguments.
  2. 2.
    A new method to dynamically integrate external knowledge retrieved from the web in LLM-based counter-argument generation.
  3. 3.
@@
    Experimental results show that the generation of counter-arguments with LLMs is improved through the integration of dynamic external knowledge, with factual evidence demonstrating a particularly significant impact on pragmatic aspects, including relatedness, persuasiveness, and factuality.
  4. 4.
    Our automatic evaluation based on LLM-as-a-Judge reveals a higher correlation with human judgments compared with reference-based metrics such as BLEU, METEOR, or BERTscore.
@@ 2 Related Work
Prior research in efficient and persuasive argument generation has approached the problem from various perspectives, focusing on different aspects of arguments. For instance, Jo et al. (2020) investigated the ability of machine learning approaches to detect attackable sentences in the arguments, and they concluded that automatic approaches are more stable in this task than human annotators, depending on the subjectivity, topic, and tone of the argument. Alshomary et al. (2021) focused on generating counter-arguments by pinpointing and challenging weak premises. Schiller et al. (2021) produces arguments by specifying the desired aspect and stance in a sentence-level setting via Conditional Transformer Language Model (CTRL) (Keskar et al., 2019). Similarly, Saha and Srihari (2023) proposed a method to control both the topic and stance of an argument while enriching it with factual evidence using an encoder-decoder language model. Alshomary and Wachsmuth (2023) propose a counter-argument generation based on the stance of the conclusion of the argument. Lin et al. (2023) employed large language models for sentence-level counter-argument generation, implementing a Chain-of-Thought methodology. Hu et al. (2023) introduces an agent-based chain-of-thought (CoT) method for generating arguments by decomposing the claim and sequentially processing the argument for the final output. Finally, Chen et al. (2024) evaluated LLMs in several argument mining and generation tasks, showing the potential of LLMs for this particular task.
The use of external sources has demonstrated its effectiveness in the generation of alternative perspectives. Wachsmuth et al. (2018) analyzed the question of retrieving the best counter-arguments when no prior knowledge is available. Hua and Wang (2018) introduced a framework that incorporated information retrieval, but their approach was limited to using the Wikipedia database as the external source. However, Wikipedia primarily contains static factual information, which may not align with the dynamic nature of arguments. To address this limitation, Stab et al. (2018) expanded the scope by indexing all documents from the Common Crawl database for argument retrieval. Hua et al. (2019) proposed an enhanced framework that leverages databases from news outlets alongside Wikipedia to retrieve evidence and improve the quality of the generated counter-arguments.
@@ 3 Data
In order to perform our experiments and to guarantee a balance in the length of input argument and output counter-argument in the data, we constructed a new corpus for the evaluation of counter-argument generation. Previous work often focused on either (too short) sentence-level (Lin et al., 2023; Schiller et al., 2021) or (too long) paragraph-level (Hua et al., 2019; Hua and Wang, 2018) generation.
Given the capabilities of modern LLMs, it was essential to specify clear input and output data to ensure accurate, robust, and fair comparisons during evaluation. Thus, our objective is to create a dataset of argument/counter-argument pairs that would meet specific criteria. We argue that (i) generating single-sentence claim-based counter-arguments is insufficient to accurately assess the quality of counter-arguments produced by LLMs and (ii), counter-arguments that are too long make it extremely difficult to properly evaluate the argumentative quality of the generated text. Our analysis revealed that Large Language Models (LLMs), when not given explicit length constraints, tend to generate verbose, essay-like responses that frequently deviate from true argumentative form, lack substantiating evidence, and demonstrate poor coherence with the original argument while overemphasizing persuasive elements. To address these limitations, we propose generating counter-arguments with a maximum length of three sentences, focusing on conciseness, factual content, and direct alignment with the input argument.
With this aim in mind, we constructed a new dataset of argument and counter-argument pairs using the CANDELA corpus as a basis Hua et al. (2019). The corpus consists of debates and discussions on various controversial topics from the r/ChangeMyView [^2] https://www.reddit.com/r/changemyview subreddit. The corpus is centred around real-world online debates where users post their opinions and evidence for any controversial topic and expect other users to provide reasoning for an alternative perspective.
@@
Original 16 372
Intermediate 3 83
Final 3 61
@@
Original 30 921
Intermediate 5 165
Final 3 72
@@
CANDELA is available in a format split by sentences, tokenized, and lowercased, making the reconstruction of the corpus necessary. To address this, we employed LLMs to convert the data back into a coherent, human-readable format.
Once the data was fully reconstructed in a human-readable format, we summarized all arguments and counter-arguments. To avoid any bias, we choose a language model different from those used in our experimentation, namely, Llama-3.1-70B-Instruct Dubey et al. (2024). The generated summaries were then manually verified against the original data and re-summarized when required to ensure semantic and pragmatic correctness.
While the corpus includes data from real-world interactions and reflects arguments from natural exchanges of information, we found that not all topics were equally suitable to evaluate whether external knowledge benefits counter-argument generation. Specifically, topics lacking a scientific or factual foundation, namely, expressed on deeply subjective topics, often trigger LLMs to produce generic, safe responses due to their safety guardrails. Moreover, the external knowledge of the topics that lack ground-truth factual backing tends to be ambiguous and/or biased. An example of such arguments can be found in Appendix B.
@@
Finally, we further manually refined the summaries to follow a structured argumentative format, emphasizing components such as the main claim, supporting evidence, and examples where applicable (see examples in Appendix C). This process resulted in a dataset of 150 high-quality 3-sentence paragraph-level argument and counter-argument pairs.
Examples of the final version of the corpus used in the experiments can be seen in Appendices A and C. The average distribution of sentences and tokens in the data is shown in Table 1.
@@ 4 Experimental setup
External knowledge is obtained through the web search tool provided by the Cohere API [^3] https://cohere.com/. The process involved automatically generating five queries (averaging 67 words each) designed to challenge the validity and veracity of the original argument’s claims and premises by questioning key points that required factual substantiation. These queries were sequentially submitted to the web search engine, and the retrieved results, averaging 5,496 words in length, were incorporated as contextual information in the final prompt presented to the language model. The model then generated counter-arguments based on both the original argument and the retrieved contextual information.
@@
The experimentation is performed using two LLMs with strong results on text generation tasks in two different configurations: (i) the LLM using only the claim as a prompt to rely on its parametric knowledge and (ii) the LLM with dynamic external knowledge. This results in the following four models:
@@
Importantly, all experiments are conducted in an inferential setting under default hyperparameters to assess the real capabilities of LLMs in counter-argument generation. This setup ensures a fair and robust evaluation of their performance in generating meaningful, well-reasoned, and factual counter-arguments. A list of all the prompts used at each generation step is illustrated in Appendix D.
@@ 4.1 Evaluation
It should be noted that the evaluation of generated counter-arguments is inherently subjective. While human evaluation is the gold standard, it is time-intensive, costly, and prone to individual biases Wachsmuth et al. (2017); Hua et al. (2019); Chen et al. (2024). To address these limitations, along with human evaluation, we will also provide two types of automatic evaluation. First, using reference-based metrics such as BLEU, METEOR, or BERTscore previously used in counter-argumentation generation Alshomary et al. (2021); Chen et al. (2024). Second, we propose to use the LLM-as-a-Judge approach. Leveraging the consistency, scalability, and efficiency of LLMs, this method enables rapid and reproducible scoring across the five dimensions. To the best of our knowledge, we are the first to propose an LLM-based evaluation for counter-argument generation.
@@
Human evaluation. We recruited human evaluators through the crowdsourcing platform Prolific [^4] https://www.prolific.com/ to assess the quality of a sample of 75 generated counter-arguments across the five predefined dimensions. We performed two rounds of evaluations of three participants to ensure that the obtained manual judgments were of high quality. We set test questions in the questionnaire to determine whether the evaluations were performed fairly. Subsequently, we selected the four participants who correctly passed the test. All evaluators were compensated above the minimum rate recommended by the platform. An example of instructions provided for manual evaluation can be found in Appendix F.
@@
Reference-based evaluation. Following the previous evaluations on counter-argument generation Hua et al. (2019); Lin et al. (2023); Chen et al. (2024) we evaluate using reference-based overlap and similarity metrics, such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), METEOR (Banerjee and Lavie, 2005), and BERTScore (Zhang et al., 2019). BLEU and ROUGE compare the overlap between the counter-argument and the claim, whereas METEOR also considers synonyms, paraphrases, and word stems. Finally, BERTScore takes into account the semantic context and meaning of the text, going beyond surface-level word matching.
Ranking. Based on the evaluation scores per dimension, we obtain rankings of each of the 5 candidate counter-arguments (4 automatically generated and the gold reference) by summing the scores for each counter-argument, as provided by the human and LLM-as-a-Judge evaluators. Formally, let the evaluation process involve $n$ counter-arguments, $m$ dimensions, and $e$ evaluators. Let $Score_{i,j,d}$ represent the score assigned by evaluator $j$ for counter-argument $i$ on dimension $d$.